OpenStudy (anonymous):

y=1/2x - 3 in standard form using integers?

OpenStudy (texaschic101):

standard form : Ax + By = C first you need to multiply everything by 2 to get rid of the fraction because an integer is a whole number, not a fraction. y = 1/2x - 3 (2)y = (2)1/2x - (2)3 2y = x - 6 Now you want to move the x to the other side and to do this you need to subtract x from both sides. 2y - x = x - x - 6 (combine like terms) 2y - x = 6 you want it in standard form Ax + By = C -x + 2y = 6 You do not want the x to be a negative so multiply the equation by a -1 (-1)-x + (-1)2y = (-1)6 x + (- 2y) = -6 <----standard form